OpenStudy (anonymous):

use factoring to solve y= x^2+2x-24

OpenStudy (anonymous):

6 & -4

OpenStudy (anonymous):

since it starts with x^2, the first part of each binomial will start with x. Since the second +/- symbol is negative, the signs in each binomial will be opposite. (x+ )(x- ) Now we have to fill in the blanks with 2 numbers that multiply to give us 24 but subtract to give us 2 (the middle term). y=(x+6)(x-4) if you are trying to find the zeros, set each binomial equal to 0 and solve for x. 0=x+6, x=-6 0=x-4, x=4 Hope this helps

OpenStudy (anonymous):

It does but im confused on which number comes first b/c there is a (4,-6) and (6,-4) to choose from

OpenStudy (radar):

You must pay attention to both the sign and magnitude. The middle term is Positive, positive 2x. You must put the one first that when combined results in a positive 2x. in this case the 6 must be positive so that the results or positive. If you used a -6 then you would of had a -2x which is not what you wanted.

OpenStudy (anonymous):

oooo ok